### v1

# Geodesic completeness, curvature singularities and infinite tidal forces

Abstract
We report some new findings regarding the subtle relations among geodesic completeness, curvature singularities and tidal forces. It is well known that any particle may be torn up near a singularity at the center of a black hole due to the divergent tidal force. However, we find that singularity is not the only cause of tidal force divergence. Even on the surface of the Earth, the tidal force experienced by a particle could be arbitrarily large if the particle moves arbitrarily close to the speed of light in a nonradial direction. Usually, a curvature singularity implies geodesic incompleteness since in many cases the metric cannot be defined at the location of the singularity. Counterexamples have been found in recent years, suggesting that geodesics could pass through curvature singularities. By taking into account the fact that any real particle is an extended body, we calculate the tidal force acting on the particle in a static and spherically symmetric spacetime. We explicitly show that an infinite tidal force always occurs near such a singularity. Therefore, no particle can actually reach the curvature singularity even if the metric is well defined at that point. We also show that the tidal acceleration along a null geodesic at the coordinate origin is divergent. For a wormhole solution that has a curvature singularity at the throat but possesses a regular metric, we show that the tidal force on any particle passing through the throat is divergent.
1 Introduction
Singularity is a fundamental prediction of general relativity, with their existence rigorously established by classical singularity theorems [1][3]. On the other hand, singularities inevitably lead to pathological behaviors of spacetimes [7]. Efforts to resolve singularity problems have been made by circumventing some of the assumptions in the singularity theorems[4] or exploring quantum effects to avoid singularities [5][6].
In fact, there are different ways to define singularities[7]. The most common ones include the curvature singularity and geodesic incompleteness. Curvature singularity usually means that the curvature invariants are divergent somewhere. Among all the invariants, seventeen curvature invariants, called Zakhary-Mcintosh (ZM) invariants, form a complete set[8]-[11]. For spherically symmetric spacetimes, it has been shown [12] that the Kretschmann scalar $K=R_{abcd}R^{abcd}$ and the Ricci square $R_{ab}R^{ab}$ are sufficient to determine whether all the seventeen ZM invariants are finite or infinite. Geodesic incompleteness refer to that the affine parameter along a timelike or null geodesic can not be extended to arbitrarily large values in either the future or past direction.
The two definitions are consistent in many cases, such as the central singularities in the Schwarzschild black hole and in the Kerr black hole. However, counterexamples have been found. For the well known C-metric[13]-[15], there exists a conic singularity while all polynomial curvature scalars are finite. On the other hand, spacetimes with curvature singularities are found to be geodesically complete [16, 17]. This may seem incredible since metric usually cannot exist where a curvature singularity appears. In fact, a regular metric can generate a spacetime with curvature singularities and a series of solutions have been explicitly constructed [17]. Unlike the usual curvature singularity where the metric cannot be defined, the metric is well defined where the curvature blows up. In such spacetimes, particles or light can reach or pass through the singularity. These spacetimes should be distinguished from regular black holes where the metric is regular everywhere and there are no central singularities. [18]-[23].
Despite the controversial definitions of singularities, it is of interest to explore their physical implications. If a particle passes through a curvature singularity, it is natural to ask whether observational effects would arise. Any real particle is an extended body, rather than a point. It is well known that extended bodies in gravitational field will experience tidal forces . The tidal force is an effect of spacetime curvature and plays an important role in astrophysics. It is closely related to gravitational waves[24]. Meanwhile, a star may be disrupted due to tidal forces produced by a black hole[25], called the tidal disruption event (TDE). TDEs may power bright flares as ultraviolet[26] and optical[27] radiation. In addition, analyses suggest that tidal accelerations can describe relativistic flow formation[28]-[30]. In previous works, tidal effects have been studied for numerous exact solutions to Einstein’s equations, including the Schwarzschild spacetime[31]-[33], the Reissner-Nordström spacetime[34][35], the Kerr spacetime[36]-[38], and regular black holes [39][40].
In this paper, we reveal that the relationships among regular metrics, curvature singularities, and tidal accelerations are far more complex and subtle. It is commonly believed that infinite curvature would result in infinite tidal forces and vice versa. In Section 2, we explicitly demonstrate that tidal acceleration can become infinite even on the Earth’s surface, provided the particle approaches the speed of light very closely. This is essentially the “Lorentz boost ” effect in gravitational field, which was previously studied in [41] [42], though the tidal effects were only implicitly discussed. We demonstrate that an extended particle moving tangentially will experience infinite tidal forces in the radial direction when its speed relative to static observers is arbitrarily close to the speed of light. In Section 3, we compute the tidal acceleration in a spherically symmetric spacetime that is geodesically complete but contains a central singularity. We show that the tidal acceleration is divergent at the central singularity, indicating that the curvature singularity may prevent the particle from reaching the origin. In addition, we find that the tidal acceleration of null geodesics similarly diverges at the origin. In Section 4, we calculate the tidal acceleration for a wormhole solution which has a curvature singularity at its throat. Although a timelike geodesic could pass through the wormhole, the tidal acceleration is always divergent at the throat, implying that the particle would be destroyed before reaching the throat. Concluding remarks are made in Section 5.
2 Tidal force on the Earth’s surface
Consider a static, spherically symmetric spacetime described by the metric
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}\,.$ (1)
The static observer has the four-velocity
$U^{a}=\frac{1}{\sqrt{f(r)}}\left(\frac{\partial}{\partial t}\right)^{a}\,.$ (2)
The tidal acceleration of a test particle is given by
$A^{a}=Z^{c}\nabla_{c}(Z^{b}\nabla_{b}W^{a})\,,$ (3)
where $Z^{a}$ is represents the four-velocity of the particle and $W^{a}$ is the deviation vector satisfying the Lie transport condition ${\mathcal{L}}_{Z}W^{a}=0$. From Eq. (3), we may obtain the familiar expression [7]
$A^{a}=R_{bcd}\hskip 0.28453pt^{a}W^{b}Z^{c}Z^{d}\,,$ (4)
where $R_{bcd}\hskip 0.28453pt^{a}$ is the Riemann tensor.
From Eq. (4), we observe that extreme tidal acceleration can arise not only from huge components of the curvature but also from huge components of $Z^{a}$. This happens when the relative speed with respect to the static observer is near the speed of light. In the following, we shall show explicitly that tangential motions could result in infinite tidal accelerations while radial motions could not.
2.1 Tangential motion
Now consider that the particle moves only in the $\phi$ direction with the four-velocity
$Z^{a}=\frac{\gamma}{\sqrt{f(r)}}\left(\frac{\partial}{\partial t}\right)^{a}+\frac{\sqrt{\gamma^{2}-1}}{r\sin\theta}\left(\frac{\partial}{\partial\phi}\right)^{a}\,.$ (5)
Note that $Z^{a}U_{a}=-\gamma$. Thus, $\gamma$ is just the familiar quantity $\gamma=\frac{1}{\sqrt{1-v^{2}}}$, where $v$ is the 3-velocity measured by the static observer.
We choose the deviation vector as
$W^{a}=\sqrt{f(r)}\left(\frac{\partial}{\partial r}\right)^{a}\,.$ (6)
It is easy to check that $Z^{a}W_{a}=0$, $Z^{a}Z_{a}=-1$ and $W^{a}W_{a}=1$. Then straightforward calculation shows
$|A|^{2}\equiv g_{ab}A^{a}A^{b}=\frac{\left[\left(\gamma^{2}-1\right)f^{\prime}(r)-\gamma^{2}rf^{\prime\prime}(r)\right]^{2}}{4r^{2}}\,.$ (7)
For the Schwarzschild solution $f(r)=1-\frac{2M}{r}$, we have
$|A|^{2}=\frac{M^{2}(1-3\gamma^{2})^{2}}{r^{6}}\,.$ (8)
Thus, when $\gamma\to\infty$, i.e., the particle moves arbitrarily close to the speed of light and the tidal acceleration is divergent.
We now quantitatively estimate the tidal acceleration experienced by a proton accelerated to ultrarelativistic velocities near the Earth’s surface. In the high Lorentz factor regime ($\gamma\gg 1$), the magnitude of the tidal acceleration in SI units scales as
$|A|\sim G\frac{M\gamma^{2}}{r^{3}}\,.$ (9)
This is the tidal acceleration per unit length. The tidal force on a proton is then estimated by
$F\sim m_{p}|A|r_{p}=G\frac{M\gamma^{2}}{r^{3}}m_{p}r_{p}\,,$ (10)
where $m_{p}$ and $r_{p}$ are the mass and radius of proton, respectively. In the Large Hadron Collider (LHC), protons can reach $99.999999\%$ of the speed of light, corresponding to $\gamma\sim 7000$. Substituting the values $m_{p}\sim 10^{-27}kg$, $r_{p}\sim 10^{-15}m$, the mass of the earth $M\sim 10^{24}kg$, the radius of the earth $r\sim 10^{6}m$, we have
$F\sim 10^{-40}N\,.$ (11)
According to [43], the local force acting on the struck quark in a proton is on the order of $3GeV/fm\sim 10^{5}N$. Although the force is amplified by a factor of $\gamma^{2}\sim 10^{7}$ due to the Lorentz boost effect, it is remains negligible and is far from being sufficient to disrupt the proton. However, particles with much higher energies have been found in cosmic rays. We expect that tidal forces on elementary particles may have detectable effects in future experiments.
2.2 Radial motion
Assume that the particle moves in the radial direction with the four velocity
$Z^{a}=\frac{\gamma}{\sqrt{f(r)}}\left(\frac{\partial}{\partial t}\right)^{a}+\sqrt{\gamma^{2}-1}\sqrt{f(r)}\left(\frac{\partial}{\partial r}\right)^{a}\,.$ (12)
Again, we have $U^{a}Z_{a}=-\gamma$.
We now consider a general deviation vector $W^{a}$ in the form
$W^{a}=W_{t}\left(\frac{\partial}{\partial t}\right)^{a}+W_{r}\left(\frac{\partial}{\partial r}\right)^{a}+W_{\theta}\left(\frac{\partial}{\partial\theta}\right)^{a}\,.$ (13)
By requiring $Z^{a}W_{a}=0$ and $W^{a}W_{a}=1$, we have
$W_{r}$ $=$ $\frac{\gamma f}{\sqrt{\gamma^{2}-1}}W_{t}$ (14)
$W_{\theta}$ $=$ $\frac{\sqrt{\gamma^{2}-1-W_{t}^{2}f}}{r\sqrt{\gamma^{2}-1}}$ (15)
Then the magnitude of the tidal acceleration is
$|A|^{2}\equiv g_{ab}A^{a}A^{b}=\frac{r^{2}W_{t}^{2}f(r)f^{\prime\prime}(r)^{2}+f^{\prime}(r)^{2}\left(\gamma^{2}-W_{t}^{2}f(r)-1\right)}{4\left(\gamma^{2}-1\right)r^{2}}\,.$ (16)
Eq. (15) shows $W_{t}^{2}f<\gamma^{2}-1$. Thus,
$|A|^{2}<\frac{r^{2}f^{\prime\prime 2}(\gamma^{2}-1)+f^{\prime 2}(\gamma^{2}-1)}{4r^{2}(\gamma^{2}-1)}=\frac{r^{2}f^{\prime\prime 2}+f^{\prime 2}}{4r^{2}}\,.$ (17)
Therefore, $|A|^{2}$ is not divergent as $\gamma\to\infty$. This indicates that an ultrarelativistic particle moving in the radial direction will not experience infinite tidal forces.
3 Tidal force at the center of a spherically symmetric spacetime
3.1 Curvature singularity and tidal force
In the previous section, we have shown that ultrarelativistic particles can experience extremely high tidal forces even in the presence of a weak gravitational field. In this section, we shall investigate an alternative mechanism that gives rise to divergent tidal forces. This mechanism relies on curvature singularity, instead of the particle’s speed.
The Computation of the Kretschmann curvature from metric (1) yields
$K=R_{abcd}R^{abcd}=\frac{4-8f+4f^{2}+4r^{2}f^{\prime 2}+r^{4}f^{\prime\prime 2}}{r^{4}}\,,$ (18)
We are interested in a regular metric at the coordinate origin $r=0$ that may be assumed in the form
$f(r)=1+C_{1}r+C_{2}r^{2}+C_{3}r^{3}+\mathcal{O}\left(r^{4}\right)\,.$ (19)
Note that the leading term must be exactly unity ($1$) to avoid introducing a conical singularity at the origin ($r=0$). Then Eq. (18) becomes
$K=\frac{8C_{1}^{2}}{r^{2}}+\frac{24C_{1}C_{2}}{r}+\mathcal{O}\left(r^{0}\right)\,.$ (20)
Thus, the necessary and sufficient condition for $K$ being divergent is $C_{1}\neq 0$. In this case, the metric is regular at $r=0$ while the curvature is singular.
We shall check if a particle passing through the origin can suffer infinite tidal acceleration. Without loss of generality, we consider that the particle moves in the equatorial plane $\theta=\pi/2$ with the four-velocity
$Z^{a}=\dot{t}\left(\frac{\partial}{\partial t}\right)^{a}+\dot{r}\left(\frac{\partial}{\partial r}\right)^{a}+\dot{\phi}\left(\frac{\partial}{\partial\phi}\right)^{a}\,,$ (21)
where the “dot” represents the derivative with respect to the proper time of the particle. The conservation of energy and angular momentum yields
$E$ $=$ $-Z_{a}\left(\frac{\partial}{\partial t}\right)^{a}\,,$ (22)
$L$ $=$ $Z_{a}\left(\frac{\partial}{\partial\phi}\right)^{a}\,.$ (23)
Along with the normalization condition $Z^{a}Z_{a}=-1$, we can solve for the components of $Z^{a}$ as
$\dot{t}=\frac{E}{f(r)},$ (24)
$\dot{\phi}=\frac{L}{r^{2}},$ (25)
$\dot{r}^{2}=E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right).$ (26)
We choose the deviation vector to be
$W^{a}=\frac{1}{r}\left(\frac{\partial}{\partial\theta}\right)^{a}\,,$ (27)
which satisfies $W^{a}W_{a}=1$ and $W^{a}Z_{a}=0$. It is then straightforward to calculate the magnitude of the tidal acceleration as
$|A|=\frac{2L^{2}[1-f(r)]+(r^{3}+L^{2}r)f^{\prime}(r)}{2r^{4}}\,.$ (28)
We first consider the radial motion, i.e., $L=0$. Straightforward computation yields
$|A|=\frac{r^{3}f^{\prime}(r)}{2r^{4}}\,,$ (29)
For the regular metric given in Eq. (19), we find
$|A|=\frac{C_{1}}{2r}+C_{2}+\frac{3C_{3}r}{2}+O(r^{2})\,.$ (30)
If the Kretschmann curvature $K$ is finite, i.e., $C_{1}=0$ as we have found above, we see that $|A|$ is finite at the origin.
If $C_{1}\neq 0$, it is easy to see that both $K$ and $|A|$ are divergent at the origin. Thus, for a particle moving in the radial direction, a divergent $K$ at the origin always implies a divergent $|A|$.
Now we turn to non-radial motions with $L\neq 0$. Define the effective potential
$V(r)=f(r)\left(1+\frac{L^{2}}{r^{2}}\right)\,.$ (31)
Eq. (26) becomes
$\dot{r}^{2}=E^{2}-V(r)\,.$ (32)
Since $f(r)>0$, we see that the particle cannot reach the origin due to the infinite potential at $r=0$. However, if $L$ is chosen to be sufficiently small, the particle can be sufficiently close to the origin. Since we are interested in the region around the origin, it is enough to take the approximation $f\approx 1$. Denote $\epsilon$ the minimum distance to the origin that the particle can reach. Noting that $\dot{r}=0$ when $r=\epsilon$, we have
$\epsilon=\sqrt{\frac{1}{E^{2}-1}}L\,.$ (33)
For $C_{1}=0$, i.e., $K$ is finite, Eq. (28) becomes
$|A|=\frac{C_{3}L^{2}}{2r}+C_{2}+\frac{3C_{3}r}{2}+O(r^{2})\,.$ (34)
By replacing $r$ in Eq. (34) with $\epsilon$ in Eq. (33), we see immediately that $|A|$ is finite.
For $C_{1}\neq 0$ ($K$ is divergent at $r=0$), we have
$|A|=-\frac{C_{1}L^{2}}{2r^{3}}+C_{2}+O(r)\,.$ (35)
Since $\epsilon\sim L$ by Eq. (33), $|A|$ is infinite at $r=0$ as long as $K$ is infinite.
In summary, a finite Kretschmann scalar $K$ ensures that the tidal acceleration remains finite at the origin. Conversely, a divergent $K$ indicates that the tidal acceleration also is infinite.
3.2 An example
The regular black hole soltuion with a nonlinear electrodynamics source, as analyzed in [17], is described by the metric function
$f(r)=1-\frac{2mr}{q^{2}+r^{2}}\,,$ (36)
where $m$ and $q$ represent the black hole’s mass and electric charge, respectively. This metric function ensures both regularity at $r=0$ and asymptotic flatness at infinity. Through Taylor expansion about $r=0$, we have
$f(r\sim 0)=1-\frac{2mr}{q^{2}}+\frac{2mr^{3}}{q^{4}}+\mathcal{O}(r^{5})\,.$ (37)
Obviously, this metric is regular at $r=0$ with $C_{1}=-\frac{2m}{q^{2}}$. The Kretschmann curvature is
$K=\frac{16m^{2}\left(16m^{4}r^{4}-32m^{3}q^{2}r^{3}+29m^{2}q^{4}r^{2}-12mq^{6}r+2q^{8}\right)}{r^{2}\left(q^{2}-2mr\right)^{6}}\,,$ (38)
which is divergent at $r=0$ as expected.
We first consider the radial motion ($L=0$). Eq. (26) becomes
$\dot{r}^{2}=E^{2}-\left(1-\frac{2mr}{q^{2}+r^{2}}\right)\,.$ (39)
Thus, the particle can reach and pass through the origin. According to Eq. (30), the tidal acceleration diverges at the origin as
$|A|\sim\frac{C_{1}}{2r}=-\frac{m}{q^{2}r}\,.$ (40)
For non-radial motion, the effective potential in Eq. (31) is described by Fig. 1. It shows that an infinite potential barrier always exists at $r=0$ and then no particle can reach it. According to Eqs. (33) and (35), $|A|$ diverges as $1/\epsilon$.
Figure 1: For nonradial motion, the potential $V(r)$ goes to infinity at $r=0$.
3.3 Null geodesics
Now we examine geodesic deviation of null geodesics. Let $k^{a}$ be the tangent of a null geodesic at the equatorial plane. By replacing $Z^{a}$ with $k^{a}$ in Eqs. (22) and (23), we can define the conserved energy and angular momentum for the null geodesic. Together with the null condition $k^{a}k_{a}=0$, we have
$k^{a}=\frac{E}{f(r)}\left(\frac{\partial}{\partial t}\right)^{a}-\sqrt{E^{2}-f(r)\frac{L^{2}}{r^{2}}}\left(\frac{\partial}{\partial r}\right)^{a}+\frac{L}{r^{2}}\left(\frac{\partial}{\partial\phi}\right)^{a}\,.$ (41)
Analogous to the timelike case, the tidal acceleration of null geodesics is defined by
$A^{a}=k^{c}\nabla_{c}(k^{b}\nabla_{b}W^{a})\,,$ (42)
where $W^{a}$ is the deviation vector satisfying ${\cal L}_{k}W^{a}=0$ and $k^{a}W_{a}=0$. It follows from Eq. (42) that
$A^{a}=R_{bcd}\hskip 0.28453pt^{a}W^{b}k^{c}k^{d}\,.$ (43)
It is evident from Eq. (42) that $A^{a}$ represents the relative acceleration of two nearby null geodesics. Unlike timelike geodesics, null geodesics exhibit purely relative acceleration without an associated “tidal force” interpretation.
Define two spacelike vector fields
$(e_{1})^{a}$ $=$ $\frac{1}{r}\left(\frac{\partial}{\partial\theta}\right)^{a},$ (44)
$(e_{2})^{a}$ $=$ $-\frac{r}{Lf(r)}\sqrt{E^{2}-f(r)\frac{L^{2}}{r^{2}}}\left(\frac{\partial}{\partial t}\right)^{a}+\frac{Er}{L}\left(\frac{\partial}{\partial r}\right)^{a}\,,$ (45)
which satisfy
$k^{a}(e_{i})_{a}=0,\qquad(e_{i})^{a}(e_{j})_{a}=\delta_{ij}\,,\qquad i,j=1,2\,.$ (46)
Thus, $\{(e_{1})^{a},(e_{2})^{a}\}$ form a two dimensional space orthogonal to $k^{a}$. A general deviation vector can be expressed as
$W^{a}=\alpha(e_{1})^{a}+\beta(e_{2})^{a}\,.$ (47)
For the regular metric in the form of Eq. (19), we have
$|A|^{2}\sim\frac{C_{1}^{2}L^{4}(\alpha^{2}+\beta^{2})}{4r^{6}}\,.$ (48)
For a radial null geodesic ($L=0$), we see immediately that $|A|=0$.
For $L\neq 0$, the radial equation
$\dot{r}^{2}=E^{2}-f(r)\frac{L^{2}}{r^{2}}$ (49)
gives the minimum distance to the origin
$\epsilon\sim L\,.$ (50)
Therefore, $|A|$ diverges near the origin as $1/\epsilon$.
4 Tidal forces near a wormhole
In contrast to conventional spherically symmetric spacetimes with a central origin at $r=0$, a wormhole geometries are characterized by a minimum surface called the throat. In this section, we explore the nature of singularity at the throat. Olmo, et at. investigated a spherically symmetric wormhole generated by the interplay between the electric field and the Palatini gravity [44]. The corresponding metric has the form
$\mathrm{d}s^{2}=-A(x)\mathrm{d}t^{2}+\frac{1}{A(x)\sigma_{+}^{2}}\mathrm{d}x^{2}+r^{2}(x)\mathrm{d}\Omega^{2},$ (51)
where
$A(x)=\frac{1}{\sigma_{+}}\left[1-\frac{r_{S}}{r}\frac{1+\delta_{1}G(r/r_{c})}{\sigma_{-}^{1/2}}\right],$ (52)
$\delta_{1}=\frac{1}{2r_{S}}\sqrt{\frac{r_{q}^{3}}{l_{\epsilon}}},\ \ \ \ \ r_{c}=\sqrt{l_{\epsilon}r_{q}},$ (53)
$\sigma_{\pm}=1\pm\frac{r_{c}^{4}}{r^{4}(x)},$ (54)
$r^{2}(x)=\frac{x^{2}+\sqrt{x^{4}+4r_{c}^{4}}}{2}.$ (55)
$r_{S}=2M_{0}$ is Schwarzschild radius, and the function $G(z)$ takes the form
$G(z)=-\frac{1}{\delta_{c}}+\frac{\sqrt{z^{4}-1}}{2}\left[f_{\frac{3}{4}}(z)+f_{\frac{7}{4}}(z)\right],$ (56)
[…]
where $\delta_{c}>0$ is a constant.
The throat of the wormhole is located at $r=r_{c}=\sqrt{l_{\epsilon}r_{q}}$, where $l_{\epsilon}$ and $r_{q}$ characterize the high-curvature corrections and the length scale associated with the electric charge, respectively. Note that $x=0$ at $r=r_{c}$.
The Krestschmann scalar is computed to be
$K$ $=$ $\frac{1}{4r^{4}}\left\{8A^{2}\left[r^{2}\left(\sigma_{-}^{\prime}\right){}^{2}+2\sigma_{-}^{2}\right]+r^{4}\left(A^{\prime}\right)^{2}\left(\sigma_{-}^{\prime}\right){}^{2}+16A\sigma_{-}\left(r^{2}A^{\prime}\sigma_{-}^{\prime}-2\right)+4r^{4}\sigma_{-}A^{\prime}A^{\prime\prime}\sigma_{-}^{\prime}\right.$ (58)
$\left.+4\sigma_{-}^{2}\left[r^{4}\left(A^{\prime\prime}\right)^{2}+4r^{2}\left(A^{\prime}\right)^{2}\right]+16\right\}\,.$
Here primes denote radial derivatives $\frac{\partial}{\partial r}$. Near the throat radius, $A(r)$ exhibits the asymptotic hehavior [16]
$\lim_{r\to r_{c}}A(x)=\frac{N_{q}}{4N_{c}}\frac{\delta_{1}-\delta_{c}}{\delta_{1}\delta_{c}}\sqrt{\frac{r_{c}}{r-r_{c}}}+\frac{N_{c}-N_{q}}{2N_{c}}+\mathcal{O}\left(\sqrt{r-r_{c}}\right),$ (59)
where $N_{q}$ is the number of charges defined by $N_{q}=q/e$ and $N_{c}$ is a constant. Substitution of Eq. (59) into Eq. (58) yields
$K=$ $\frac{\left(\delta_{c}-\delta_{1}\right)^{2}r_{S}^{2}}{4\delta_{c}^{2}r_{c}^{3}\left(r-r_{c}\right)^{3}}-\frac{5\left(\delta_{c}-\delta_{1}\right)^{2}r_{S}^{2}}{8\delta_{c}^{2}r_{c}^{4}\left(r-r_{c}\right)^{2}}+\frac{2\left(\delta_{c}-\delta_{1}\right)r_{S}\left(2\delta_{1}r_{S}-3r_{c}\right)}{3\delta_{c}r_{c}^{9/2}\left(r-r_{c}\right)^{3/2}}-\frac{11\left(\delta_{c}-\delta_{1}\right)^{2}r_{S}^{2}}{8\delta_{c}^{2}r_{c}^{5}\left(r-r_{c}\right)}$
$+\frac{\left(\delta_{c}-\delta_{1}\right)r_{S}\left(435r_{c}-418\delta_{1}r_{S}\right)}{30\delta_{c}r_{c}^{11/2}\sqrt{r-r_{c}}}+\mathcal{O}\left[\left(r-r_{c}\right)^{0}\right].$ (60)
Hence, there is a curvature singularity at the throat $r=r_{c}$. Nevertheless, the spacetime is geodesically complete and a timelike geodesic can pass through the throat [16]. We still consider an extended particle and calculate its tidal acceleration near the throat. In the remainder of this section, we shall only focus on the case $\delta_{1}\neq\delta_{c}$ as it corresponds to the curvature singularity.
The equations of motion of a timelike geodesic in the equatorial plane are given by
$\dot{t}=\frac{E}{A},$ (61)
$\dot{\phi}=\frac{L}{r^{2}},$ (62)
$\dot{x}^{2}=\sigma_{+}^{2}\left[E^{2}-A\left(1+\frac{L^{2}}{r^{2}}\right)\right]=\sigma_{+}^{2}\left[E^{2}-V(r)\right].$ (63)
where
$V(r)=A\left(1+\frac{L^{2}}{r^{2}}\right)=\frac{\left(\delta_{1}-\delta_{c}\right)r_{S}\left(r_{c}^{2}+L^{2}\right)}{4\delta_{c}r_{c}^{5/2}\sqrt{r-r_{c}}}+\mathcal{O}\left[\left(r-r_{c}\right)^{0}\right].$ (64)
For $\delta_{1}<\delta_{c}$, $V(r)<0$ and then there is an infinitely attractive potential at $x=0$. Consequently, particles can easily traverse the throat.
For $\delta_{1}>\delta_{c}$, the potential becomes infinitely repulsive at the throat, preventing particles from traversing it. Instead, an incoming particle is deflected and orbits the wormhole at a minimum radius $r_{m}(>r_{c})$. Since $\dot{x}$ vanishes at the turning point $r=r_{m}$, we can combine Eqs. (63) and (64) to obtain
$E^{2}=V(r_{m})=A|_{r=r_{m}}\left(1+\frac{L^{2}}{r_{m}^{2}}\right)\sim\frac{1}{\sqrt{r_{m}-r_{c}}}+\mathcal{O}\left[\left(r_{m}-r_{c}\right)^{0}\right]\,,$ (65)
which shows that the $r_{m}-r_{c}\sim E^{-4}$. Thus the particle cannot get arbitrarily close to the throat since the energy $E$ is finite.
Our goal is to determine whether the tidal acceleration is infinite near the throat. The four-velocity of a timelike particle is given by
$Z^{a}=\frac{E}{A}\left(\frac{\partial}{\partial t}\right)^{a}-\sigma_{+}\sqrt{E^{2}-V}\left(\frac{\partial}{\partial x}\right)^{a}+\frac{L}{r^{2}}\left(\frac{\partial}{\partial\phi}\right)^{a}\,.$ (66)
We choose three spacelike vectors orthogonal to $Z^{a}$ as independent deviation vectors, that are defined by
$W_{1}^{a}=-\sqrt{\frac{E^{2}-V}{VA}}\left(\frac{\partial}{\partial t}\right)^{a}+E\sigma_{+}\sqrt{\frac{A}{V}}\left(\frac{\partial}{\partial x}\right)^{a},$ (67)
$W_{2}^{a}=\frac{1}{r}\left(\frac{\partial}{\partial\theta}\right)^{a},$ (68)
$W_{3}^{a}=\sqrt{\frac{L^{2}}{L^{2}+r^{2}}}\left[\frac{E}{A}\left(\frac{\partial}{\partial t}\right)^{a}-\sigma_{+}\sqrt{E^{2}-V}\left(\frac{\partial}{\partial x}\right)^{a}\right]+\sqrt{1+\frac{L^{2}}{r^{2}}}\frac{1}{r}\left(\frac{\partial}{\partial\phi}\right)^{a}.$ (69)
Let $A_{i}^{a}$ ($i=1,2,3$) denote the tidal acceleration corresponding to each deviation vector $W_{i}^{a}$, i.e., $A_{i}^{a}=R_{bcd}^{\ \ \ a}\left(W_{i}\right)^{b}Z^{c}Z^{d}$. We first analyze the case $\delta_{1}<\delta_{c}$, where, as previously discussed, the particle can reach the throat when $\delta_{1}<\delta_{c}$. The magnitudes of the tidal accelerations near the throat for each spatial direction are given by
$|A_{1}|^{2}$ $=$ $\frac{\left(\delta_{c}-\delta_{1}\right)^{2}r_{S}^{2}\left(r_{c}^{2}+L^{2}\right)^{2}}{16\delta_{c}^{2}r_{c}^{7}\left(r-r_{c}\right)^{3}}+\mathcal{O}\left[\left(r-r_{c}\right)^{-2}\right],$ (70)
$|A_{2}|^{2}$ $=$ $\frac{\left(\delta_{c}-\delta_{1}\right)r_{S}\left(r_{c}^{2}+L^{2}\right)}{4\delta_{c}r_{c}^{9/2}\sqrt{r-r_{c}}}+\mathcal{O}\left[\left(r-r_{c}\right)^{0}\right],$ (71)
$|A_{3}|^{2}$ $=$ $\frac{\left(\delta_{c}-\delta_{1}\right){}^{2}r_{S}^{2}}{16\delta_{c}^{2}r_{c}^{5}\left(r-r_{c}\right)}+\mathcal{O}\left[\left(r-r_{c}\right)^{-1/2}\right].$ (72)
Clearly, the three tidal accelerations associated with the three spatial directions all diverges as $r\to r_{c}$. In the case $\delta_{1}>\delta_{c}$, we have shown that the particle with finite energy cannot approach arbitrarily close to the throat. Consequently, its tidal acceleration remains finite for any deviation vector. We therefore arrive at the following conclusion: Although a particle can approach the throat of the wormhole when $\delta_{1}<\delta_{c}$, the tidal force acting on it becomes infinitely large due to the curvature singularity. This implies that the wormhole is not traversable by timelike particles.
5 Concluding remarks
In this paper, we have explored the relations among geodesic completeness, curvature singularities and tidal forces. Our key finds are as follows.
We demonstrated that the tidal force experienced by a particle could be extremely large even on the Earth. To achieve this, the particle must move non-radially at ultrarelativistic speed. This result illustrates that ultrarelativistic particles can amplify the magnitude of gravitational field. It may also indicate that the strength of gravitational field is relative and observer dependent.
For a class of spherically symmetric spacetimes with metric regular at the origin but harboring curvature singularities, we showed that a particle moving in the radial direction can pass through the origin and endure infinite tidal forces. For non-radial motions, by fine-tuning the angular momentum to arbitrarily small values, particles can approach arbitrarily close to the origin, where the tidal forces likewise diverge. If a photon travels strictly in the radial direction, its tidal acceleration associated with any deviation vector vanishes at the origin. Otherwise, the tidal acceleration for a photon diverges near the origin.
Finally, we calculated the tidal acceleration around a wormhole that is geodesically complete but possesses a curvature singularity at the throat. We found that the particle always experiences infinitely large tidal forces as it traverses the throat. This divergence renders the wormhole effectively non-traversable for timelike particles.
Our analysis demonstrates that curvature singularities are generically accompanied by divergent tidal accelerations even in geodesically complete spacetimes. We have specifically examined this relationship through the Kretschmann scalar in spherically symmetric geometries. It would be worthwhile to extend the discussion to other curvature invariants in more general spacetime backgrounds.
Acknowledgement
We thank Prof. Robert Wald for valuable comments on the relationship between the tidal force and Lorentz boost as well as recommending the references [41]-[42]. We also thank Prof. Xiaoning Wu for helpful discussions.